What is the sign and magnitude of a point charge that produces a potential of -3.00 V at a distance of 7.00 mm?

____ C
Answer #1

F=\frac{KQ}{r^2}

and for potential energy

V=\frac{KQ}{r}

-3=\frac{9X10^9XQ}{7X10^{-3}m}

Q = -2.33X10-12 C